# Ensemble Kalman Filtering Meets Gaussian Process SSM for Non-Mean-Field and Online Inference

### v3

@@ Abstract
Gaussian process state-space models (GPSSMs) are a versatile and principled family of nonlinear dynamical system models. However, existing variational learning and inference methods for GPSSMs often necessitate optimizing a substantial number of variational parameters, leading to inadequate performance and efficiency. To overcome this issue, we propose incorporating the ensemble Kalman filter (EnKF), a well-established model-based filtering technique, into the variational inference framework to approximate the posterior distribution of latent states. This utilization of EnKF can effectively exploit the dependencies between latent states and GP dynamics, while eliminating the need for parameterizing the variational distribution, thereby significantly reducing the number of variational parameters. Moreover, we show that our proposed algorithm allows straightforward evaluation of an approximated evidence lower bound (ELBO) in variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging automatic differentiation tools, we hence can maximize the ELBO and train the GPSSM efficiently. We also extend the proposed algorithm to accommodate an online setting and provide detailed algorithmic analyses and insights. Extensive evaluation on diverse real and synthetic datasets demonstrates the superiority of our EnKF-aided variational inference algorithms in terms of learning and inference performance compared to existing methods.
@@ I Introduction
Gaussian processes, serving as the most prominent Bayesian nonparametric model [10], provide the flexibility to model general nonlinear system dynamics without enforcing an explicit parametric structure. With the inherent regularization imposed by the GP prior, GPSSMs are able to mitigate model overfitting and generalization issues [11], rendering them to be more effective in scenarios with limited data samples [12]. Moreover, owing to its Bayesian nature, GPSSMs maintain good interpretability and explicit uncertainty calibration for analyzing system dynamics. These superior properties have led to the extensive usage of GPSSMs in various live applications, such as human pose and motion learning [13], robotics and control learning [14], reinforcement learning [15], target tracking and navigation [16], and magnetic-field sensing [17].
Despite the popularity of GPSSMs, simultaneously learning the model and estimating the latent states in GPSSMs remains highly challenging, primarily due to the following two factors. First, the quality of the state estimates affects the model learning and vice versa, leading to heightened computational and statistical complexities. Second, the nonlinearity in GPSSMs prohibits tractable learning and inference process [10]. Hence, the main task in GPSSM is to accurately approximate the joint posterior over the latent states and the system dynamics represented by GPs. Multiple approaches have emerged towards achieving this objective in the last decade.
The seminal work utilizing particle Markov chain Monte Carlo (PMCMC) was proposed in [8]. Subsequently, more advanced methods [18, 19, 20, 21] leveraged the reduced-rank GP approximation introduced in [22] to alleviate the substantial computational demands of GPs in [8]. However, the computational burden of the involved PMCMC remains unaffordable, particularly when dealing with high-dimensional latent states. Consequently, there has been a paradigm shift towards variational inference methods [23, 24, 25, 26, 27, 28, 29, 30, 31], which adopted the classic sparse GP approximations with inducing points [32].
Generally, the variational inference approaches can be classified into two main categories: the mean-field (MF) class [23, 24, 25, 26] versus the non-mean-field (NMF) class [27, 28, 29, 30, 31], depending on whether the statistical independence assumption has to be applied to the variational distribution of the latent states and GP dynamics. The first variational algorithm, incorporating the MF assumption and integrating PF for latent state inference, was introduced in [23]. Subsequent MF endeavors [24, 25, 26] aimed at reducing computational and model complexities in [23]. Yet, the main idea of the MF class is to neglect the inherent dependencies between the latent states and the GP dynamics, potentially impairing the overall learning accuracy and yielding overconfident state estimates [33].
@@
To tackle the computational challenges mentioned above, our proposed methods incorporate the model-based filtering technique, EnKF, to approximate the joint posterior distribution. By leveraging the EnKF, our method can fully exploit the dependencies between the latent states and the GP dynamics, while reducing the parameterization of associated variational distributions, thereby alleviating model and computational complexities and accelerating algorithm convergence. This streamlined parameterization also helps the easy accommodation of our method in an online setting. The main contributions of this paper are summarized as follows.
@@
  • By leveraging EnKF, we also demonstrate a simplified evidence lower bound (ELBO) for variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging the differentiable nature of the classic EnKF alongside the off-shelf automatic differentiation tools, we can optimize the ELBO and train the GPSSM efficiently.
  • Without explicitly parameterizing the variational distribution over latent states, the proposed EnKF-aided algorithm gets readily extended to accommodate online applications. This online method relies on a more principled objective by ensuring observation fitting accuracy while incorporating model regularizations to mitigate model overfitting, rendering it superior to state-of-the-art online algorithms in learning and inference efficiency.
  • We provide detailed analysis and fresh insights for the proposed algorithms and conduct extensive experiments on various real and synthetic datasets to evaluate the performance of these algorithms from various aspects. The results demonstrate that the proposed EnKF-aided variational learning methods consistently outperform the existing state-of-the-art methods, especially in terms of learning and inference.
The remainder of this paper is organized as follows. Some preliminaries related to GPSSMs are provided in Section II. The computational challenges of the existing methods are summarized in Section III. Section IV elaborates the proposed EnKF-aided variational learning algorithm. Section V extends the proposed algorithm to accommodate online applications. The experimental results are presented in Section VI, and Section VII concludes this paper. More supportive results, proofs, etc. are relegated to the Appendix and supplementary materials.
@@ II-A Gaussian Processes (GPs)
where $\bm{K}_{\bm{X},\bm{X}}$ denotes the covariance matrix evaluated on the training input $\bm{X}$, and each entry is $[\bm{K}_{\bm{X},\bm{X}}]_{i,j}=k({\bm{x}}_{i},\bm{x}_{j})$; $\bm{K}_{\bm{x}_{*},\bm{X}}$ denotes the cross covariance matrix between the test input $\bm{x}_{*}$ and the training input $\bm{X}$; the zero-mean GP prior is assumed here and will be used in the rest of this paper if there is no further specification. Note that the posterior distribution $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$ gives not only a point estimate, i.e., the posterior mean, but also an uncertainty region of such estimate quantified by the posterior variance. It is also noteworthy that here we denote the variables in the GPR using mathematical mode italics, such as $\bm{x}_{i}$ and $y_{i}$; these variables should not be confused with the latent state ${\mathbf{x}}_{t}$ and observation ${\mathbf{y}}_{t}$ in SSM (cf. Eq. (II-B)).
@@ II-B Gaussian Process State-Space Models (GPSSMs)
[…]
where the latent states form a Markov chain. That is, for any time instance $t\!\in\!\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ is generated by conditioning on only ${\mathbf{x}}_{t}$ and the transition function $f(\cdot)\!:\!\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$. The emission is assumed to be linear with a known coefficient matrix, $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$, hence mitigating the system non-identifiability111 Nonlinear emissions can be addressed by augmenting the latent state to a higher dimension. This augmentation helps mitigate/eliminate the significant non-identifiability issues commonly encountered in GPSSMs [35, 24]. The initial state distribution $p({\mathbf{x}}_{0})$ is also known and assumed to follow a Gaussian distribution. Both the states and observations are corrupted by zero-mean Gaussian noise with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
The GPSSM incorporates a GP prior to model the transition function $f(\cdot)$ in Eq. (II-B). Specifically, Fig. 1 presents the graphical model of GPSSM, while the following equations express its mathematical representation:
$f(\cdot)\sim\mathcal{G}\mathcal{P}\left(\mu(\cdot),k(\cdot,\cdot);{{\bm{\theta}}}_{gp}\right)$ (5a)
$\mathbf{x}_{0}\sim p\left(\mathbf{x}_{0}\right)$ (5b)
${{\mathbf{f}}}_{t}=f\left(\mathbf{x}_{t-1}\right)$ (5c)
$\mathbf{x}_{t}\mid{{\mathbf{f}}}_{t}\sim\mathcal{N}\left({\mathbf{x}}_{t}\mid{{\mathbf{f}}}_{t},\mathbf{Q}\right)$ (5d)
$\mathbf{y}_{t}\mid\mathbf{x}_{t}\sim{\cal N}\left(\mathbf{y}_{t}\mid\bm{C}\mathbf{x}_{t},\mathbf{R}\right)$ (5e)
where $\mathbf{f}_{t}$ represents the GP transition function value evaluated at the previous state $\mathbf{x}_{t-1}$. Note that when dealing with a multidimensional state space ($d_{x}\!>\!1$), the transition function $f(\cdot):\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$ is represented using a multi-output GP. In this context, the $d_{x}$ functions are typically modeled with $d_{x}$ mutually independent GPs [36]. Furthermore, it is noteworthy that the GPSSM illustrated in Fig. 1 can be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state with $[{\mathbf{x}}_{t},\bm{c}_{t}]\in\mathbb{R}^{d_{x}+d_{c}}$. For the sake of brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.
@@
Given the aforementioned model descriptions, the joint density function of the GPSSM can be expressed as:
@@
where $p({\mathbf{f}}_{1:T})\!\!=\!\!p(f({\mathbf{x}}_{0:T-1}))\!\!=\!\!\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{f}}_{1:t-1},{\mathbf{x}}_{0:t-1})$ corresponds to a finite dimensional ($T$-dimensional) GP distribution [24], and we define the short-hand notations ${\vec{\mathbf{y}}}\triangleq{\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, ${\vec{\mathbf{f}}}\triangleq{\mathbf{f}}_{1:T}=\{{\mathbf{f}}_{t}\}_{t=1}^{T}$, and ${\vec{\mathbf{x}}}\triangleq{\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$. The model parameters $\bm{\theta}$ includes the noise covariance and GP hyper-parameters, i.e., $\bm{\theta}\!\!=\!\![\mathbf{Q},\mathbf{R},\bm{\theta}_{gp}]$. The challenging task in GPSSM is to learn ${{\bm{\theta}}}$, and simultaneously infer the latent states of interest, which involves the marginal distribution $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$. However, due to the nonlinearity of GP, a closed-form solution for $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is unavailable. Hence, it becomes necessary to utilize approximation methods.
@@ III Problem Statement
$\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})-\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{p({\vec{\mathbf{f}}},{\vec{\mathbf{x}}},{\vec{\mathbf{y}}})}{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\right]}_{\triangleq\mathcal{L}}=\underbrace{\operatorname{KL}\!\!\left[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})\right]}_{\geq 0}$ (7)
Detailed derivations can be found in Supplement D-A. Maximizing $\mathcal{L}$ w.r.t. $\bm{\theta}$ fine-tunes the model parameters to fit the observed data in the model learning process; while maximizing $\mathcal{L}$ w.r.t. the variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is equivalent to minimizing the KL divergence, $\operatorname{KL}[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})]$. That is, it enhances the quality of the variational distribution approximation, bringing it closer to the underlying posterior distribution, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, which corresponds to the model inference process [37]. The capacity of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ to approximate $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$ is thus of crucial importance [38].
@@
where $\{\bm{A}_{t},\bm{b}_{t},\bm{S}_{t}\}_{t=1}^{T}$ are free variational parameters. This choice, however, introduces a significant drawback: the number of variational parameters grows linearly with the length of the time series. Although this issue can be partially addressed by incorporating an inference network, such as a bidirectional recurrent neural network [26, 39], to learn the variational parameters and make the variational distribution learning with a constant model complexity, fine-tuning these variational parameters of the inference network still requires substantial effort. Consequently, despite the flexibility of the variational distribution described in Eq. (9) allowing for approximation of the true posterior distribution, its empirical performance often falls short of its theoretical expressive capacity [30]. Moreover, it is noteworthy that the inference network commonly takes the input of the observations, ${\vec{\mathbf{y}}}$, rendering the existing inference network-based methods predominantly trained in an offline manner, thus posing challenges when attempting to adapt them for online applications. To address these limitations, we step away from the heavy parameterization (e.g., the black-box inference networks) and propose a novel interpretable EnKF-aided algorithm for variational inference in GPSSMs. Our algorithm can exploit the dependencies between ${\vec{\mathbf{x}}}$ and ${\vec{\mathbf{f}}}$ and alleviate the model and computational complexities while being easily extended to an online learning algorithm.

### v5

@@ Abstract
The Gaussian process state-space models (GPSSMs) represent a versatile class of data-driven nonlinear dynamical system models. However, the presence of numerous latent variables in GPSSM incurs unresolved issues for existing variational inference approaches, particularly under the more realistic non-mean-field (NMF) assumption, including extensive training effort, compromised inference accuracy, and infeasibility for online applications, among others. In this paper, we tackle these challenges by incorporating the ensemble Kalman filter (EnKF), a well-established model-based filtering technique, into the NMF variational inference framework to approximate the posterior distribution of the latent states. This novel marriage between EnKF and GPSSM not only eliminates the need for extensive parameterization in learning variational distributions, but also enables an interpretable, closed-form approximation of the evidence lower bound (ELBO). Moreover, owing to the streamlined parameterization via the EnKF, the new GPSSM model can be easily accommodated in online learning applications. We demonstrate that the resulting EnKF-aided online algorithm embodies a principled objective function by ensuring data-fitting accuracy while incorporating model regularizations to mitigate overfitting. We also provide detailed analysis and fresh insights for the proposed algorithms. Comprehensive evaluation across diverse real and synthetic datasets corroborates the superior learning and inference performance of our EnKF-aided variational inference algorithms compared to existing methods.
@@ I Introduction
Gaussian processes, serving as the most prominent Bayesian non-parametric model [9], provide the flexibility to model general nonlinear system dynamics without enforcing an explicit parametric structure. With the inherent regularization imposed by the GP prior, GPSSMs are able to mitigate overfitting and model generalization issues [10], rendering them to be more effective in scenarios with limited data samples [11]. Moreover, owing to its Bayesian nature, GPSSMs maintain good interpretability and explicit uncertainty calibration for analyzing system dynamics. These superior properties have led to the extensive usage of GPSSMs in various live applications, such as human pose and motion learning [12], robotics and control learning [13], reinforcement learning [14], target tracking and navigation [15], and magnetic-field sensing [16].
Despite the popularity of GPSSMs, simultaneously learning the model and estimating the latent states in GPSSMs remains highly challenging, primarily due to the following two factors. First, the inference quality of the numerous latent states affects the model learning and vice versa, leading to heightened computational and statistical complexities. Second, the nonlinearity in GPSSMs prohibits tractable learning and inference processes [17, 18, 19]. Hence, the main task in GPSSM is to accurately approximate the joint posterior over the latent states and the system dynamics represented by GPs. Multiple approaches have emerged towards achieving this objective in the last decade.
The seminal work utilizing particle Markov chain Monte Carlo (PMCMC) was proposed in [8]. Subsequently, more advanced methods [20, 21, 22, 23] leveraged the reduced-rank GP approximation introduced in [24] to alleviate the substantial computational demands of GPs in [8]. However, the computational burden of the involved PMCMC remains unaffordable, particularly when dealing with long and high-dimensional latent state trajectories. Consequently, there has been a paradigm shift towards variational inference methods [25, 26, 27, 28, 29, 30, 31, 32, 33], which adopted the classic sparse GP approximations with inducing points [34].
Generally, the variational inference approaches can be classified into two main categories: the mean-field (MF) class [25, 26, 27, 28] versus the non-mean-field (NMF) class [29, 30, 31, 32, 33], depending on whether the statistical independence assumption is applied to the variational distribution of the latent states and GP dynamics. The first variational algorithm, incorporating the MF assumption and integrating PF for latent state inference, was introduced in [25]. Subsequent MF endeavors [26, 27, 28] aimed at reducing computational and model complexities in [25]. While the MF class methods may enable more manageable computations, they neglect the inherent dependencies between the latent states and the GP dynamics, potentially impairing the overall learning accuracy and yielding overconfident state estimates [35].
@@
To tackle the computational challenges mentioned above, we propose integrating the model-based filtering technique, EnKF, into the NMF variational inference framework. The primary benefit of this approach lies in the streamlined parameterization of the associated variational distributions, improving inference tractability and accelerating algorithm convergence. Additionally, this streamlined parameterization enables seamless accommodation of our method in online learning applications. The main contributions are summarized as follows.
@@
  • By leveraging EnKF, we also demonstrate an approximation of the evidence lower bound (ELBO) via simply summating multiple interpretable terms with readily available closed-form solutions. Leveraging the differentiable nature of the classic EnKF alongside the off-shelf automatic differentiation tools, we can optimize the ELBO and train the GPSSM efficiently.
  • Without explicitly parameterizing the variational distribution over latent states, the proposed EnKF-aided algorithm gets readily extended to accommodate online learning applications. The resulting online algorithm relies on a principled objective by ensuring observation fitting accuracy while incorporating model regularizations to mitigate model overfitting, rendering it superior to state-of-the-art online algorithms in learning and inference efficiency.
  • We offer in-depth analysis and novel insights into the proposed algorithms, and conduct extensive experiments on diverse real and synthetic datasets to assess their performance from various aspects. The results demonstrate that the proposed EnKF-aided variational learning algorithms consistently outperform the existing state-of-the-art methods, especially in terms of learning and inference.
The remainder of this paper is organized as follows. Some preliminaries related to GPSSMs are provided in Section II. The computational challenges of the existing methods are summarized in Section III. Section IV elaborates the proposed EnKF-aided variational learning algorithm. Section V extends the proposed algorithm to accommodate online learning applications. The experimental results are presented in Section VI, and Section VII concludes this paper. More supportive results, proofs, etc., are relegated to the Appendix and supplementary materials [37].
@@ II-A Gaussian Processes (GPs)
where $\bm{K}_{\bm{X},\bm{X}}$ denotes the covariance matrix evaluated on the training input $\bm{X}$, and each entry is $[\bm{K}_{\bm{X},\bm{X}}]_{i,j}=k({\bm{x}}_{i},\bm{x}_{j})$; $\bm{K}_{\bm{x}_{*},\bm{X}}$ denotes the cross covariance matrix evaluated on the test input $\bm{x}_{*}$ and the training input $\bm{X}$; the zero-mean GP prior is assumed here and will be used in the rest of this paper if there is no further specification. Note that the posterior distribution $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$ gives not only a point estimate, i.e., the posterior mean, but also an uncertainty region of such estimate quantified by the posterior variance. It is also noteworthy that here we denote the variables in the GPR using mathematical mode italics, such as $\bm{x}_{i}$ and $y_{i}$; these variables should not be confused with the latent state ${\mathbf{x}}_{t}$ and observation ${\mathbf{y}}_{t}$ in SSM (cf. Eq. (4)).
@@ II-B Gaussian Process State-Space Models (GPSSMs)
$\text{({Transition})}\qquad{\mathbf{x}}_{t+1}$ $=f({\mathbf{x}}_{t})+\mathbf{v}_{t},$ $\mathbf{v}_{t}\sim{\cal N}(\bm{0},\mathbf{Q}),$ (4a)
$\text{({Emission})}\qquad\quad{\mathbf{y}}_{t}$ $=\bm{C}{\mathbf{x}}_{t}+\mathbf{e}_{t},$ $\mathbf{e}_{t}\sim{\cal N}(\bm{0},\mathbf{R}),$ (4b)
where the latent states form a Markov chain. That is, for any time instance $t\!\in\!\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ is generated by conditioning on only ${\mathbf{x}}_{t}$ and the transition function $f(\cdot)\!:\!\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$. The emission is assumed to be linear with a known coefficient matrix, $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$, hence mitigating the system non-identifiability111 Nonlinear emissions can be addressed by augmenting the latent state to a higher dimension. This augmentation helps mitigate/eliminate the significant non-identifiability issues commonly encountered in GPSSMs [38, 26]. Both the states and observations are corrupted by zero-mean Gaussian noise with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
The GPSSM incorporates a GP prior to model the time-invariant transition function $f(\cdot)$ in Eq. (4). Specifically, Fig. 1 presents the graphical model of GPSSM, while the following equations express its mathematical representation:
$f(\cdot)\sim\mathcal{G}\mathcal{P}\left(\mu(\cdot),k(\cdot,\cdot);{{\bm{\theta}}}_{gp}\right),$ (5a)
$\mathbf{x}_{0}\sim p\left(\mathbf{x}_{0}\right),$ (5b)
${{\mathbf{f}}}_{t}=f\left(\mathbf{x}_{t-1}\right),$ (5c)
$\mathbf{x}_{t}\mid{{\mathbf{f}}}_{t}\sim\mathcal{N}\left({\mathbf{x}}_{t}\mid{{\mathbf{f}}}_{t},\mathbf{Q}\right),$ (5d)
$\mathbf{y}_{t}\mid\mathbf{x}_{t}\sim{\cal N}\left(\mathbf{y}_{t}\mid\bm{C}\mathbf{x}_{t},\mathbf{R}\right).$ (5e)
where $\mathbf{f}_{t}$ represents the GP transition function value evaluated at the previous state $\mathbf{x}_{t-1}$. For multidimensional state spaces ($d_{x}\!>\!1$), the transition function $f(\cdot):\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$ is represented using a multi-output GP. In this context, the $d_{x}$ functions are typically modeled with $d_{x}$ mutually independent GPs [39]. The prior distribution of the initial state, $p({\mathbf{x}}_{0})$, is assumed to be Gaussian and known for simplicity; however, it can also be learned from observed data in the absence of prior information [29]. Additionally, it is noteworthy that the GPSSM illustrated in Fig. 1 can be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state with $[{\mathbf{x}}_{t},\bm{c}_{t}]\in\mathbb{R}^{d_{x}+d_{c}}$. For the sake of brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.
Remark 1.
It is important to note the distinction between GPSSMs and the state-space representation of GPs (SSGP). The SSGP is a method that converts a GP into a linear state-space form, enabling the use of efficient inference techniques such as the Kalman filter and smoother, thereby facilitating the computationally efficient handling of large-scale problems. For more details on SSGP, we direct readers to [40, 41]. In contrast, a GPSSM utilizes the non-parametric flexibility of GPs to model the state transitions in SSMs, allowing for the capture of complex, nonlinear system dynamics. This results in a more flexible but typically computationally intensive SSM.
@@
Given the aforementioned model, see Eq. (5), the joint density function of the GPSSM can be expressed as:
@@
where $p({\mathbf{f}}_{1:T})\!\!=\!\!p(f({\mathbf{x}}_{0:T-1}))\!\!=\!\!\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{f}}_{1:t-1},{\mathbf{x}}_{0:t-1})$ corresponds to a finite dimensional ($T$-dimensional) GP distribution [26], and we define the short-hand notations ${\vec{\mathbf{y}}}\triangleq{\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, ${\vec{\mathbf{f}}}\triangleq{\mathbf{f}}_{1:T}=\{{\mathbf{f}}_{t}\}_{t=1}^{T}$, and ${\vec{\mathbf{x}}}\triangleq{\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$. The model parameters $\bm{\theta}$ includes the noise covariance and GP hyper-parameters, i.e., $\bm{\theta}\!\!=\!\!\{\mathbf{Q},\mathbf{R},\bm{\theta}_{gp}\}$. The challenging task in GPSSM is to learn ${{\bm{\theta}}}$, and simultaneously infer the latent states of interest, which involves the marginal distribution $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$. However, due to the nonlinearity of GP, a closed-form solution for $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is unavailable. Hence, it becomes necessary to utilize approximation methods.
@@ III Problem Statement
$\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})-\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{p({\vec{\mathbf{f}}},{\vec{\mathbf{x}}},{\vec{\mathbf{y}}})}{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\right]}_{\triangleq\mathcal{L}}=\underbrace{\operatorname{KL}\!\!\left[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})\right]}_{\geq 0}.$ (7)
Detailed derivations can be found in Supplement D-A. Maximizing $\mathcal{L}$ with respect to (w.r.t.) $\bm{\theta}$ fine-tunes the model parameters to fit the observed data in the model learning process; while maximizing $\mathcal{L}$ w.r.t. the variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is equivalent to minimizing the KL divergence, $\operatorname{KL}[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})]$. That is, it enhances the quality of the variational distribution approximation, bringing it closer to the underlying posterior distribution, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, which corresponds to the model inference process [42]. The capacity of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ to approximate $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$ is thus of crucial importance [43].
@@
where $\{\bm{A}_{t},\bm{b}_{t},\bm{S}_{t}\}_{t=1}^{T}$ are free variational parameters. This choice, however, introduces a significant drawback: the number of variational parameters grows linearly with the length of the time series. Although this issue can be partially addressed by incorporating an inference network, such as a bidirectional recurrent neural network [28, 44], to learn the variational parameters and make the variational distribution learning with a constant model complexity, fine-tuning these variational parameters of the inference network still requires substantial effort. Consequently, despite the flexibility of the variational distribution described in Eq. (9) allowing for approximation of the true posterior distribution, its empirical performance often falls short of its theoretical expressive capacity [32]. Moreover, it is noteworthy that the inference network commonly takes the input of the observations, ${\vec{\mathbf{y}}}$, rendering the existing inference network-based methods predominantly trained in an offline manner, thus posing challenges when attempting to adapt them for online applications.
To address these limitations, we step away from the heavy parameterization (e.g., the black-box inference networks) and propose a novel interpretable EnKF-aided algorithm for variational inference in GPSSMs. Our algorithm can exploit the dependencies between ${\vec{\mathbf{x}}}$ and ${\vec{\mathbf{f}}}$ and alleviate the model and computational complexities while being easily extended to an online learning algorithm.